The current price of a stock is $33, and the annual risk-free rate is 6%. A call option with a strike price of $31 and with 1 year until expiration has a current value of $6.22. What is the value of a put option written on the stock with the same exercise price and expiration date as the call option? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

As per put call parity
Call Price + Strike Price*e^(-rt) = Put price + Stock Price
6.22+31*e^(-6%*1)=Put Price+33

Put Price =6.22+31*e^(-6%*1)-33 =2.41